OpenStudy (spinex12):

24.6-8.47 1.01 9.9 16.13 94.27

OpenStudy (anonymous):

ummm b

OpenStudy (anonymous):

Do you know how to subtract decimal numbers?

OpenStudy (spinex12):

no

OpenStudy (spinex12):

I am in 3rd grade

OpenStudy (alexandervonhumboldt2):

b is wrong

OpenStudy (anonymous):

ok, first organize them according to the decimal place

OpenStudy (gabbyalicorn):

24.6−8.47 =24.6−8.47 =24.6+−8.47 =16.13

OpenStudy (anonymous):

Oh, very well! So you're very smart. That's why you're solving this problem. ok, ready to learn how to subtract?

OpenStudy (spinex12):

ok

OpenStudy (anonymous):

so you have 24.6 -84.7 align it by the decimal points

OpenStudy (anonymous):

Why not count up? Find the distance between 8.47 and 9, then the distance between 9 and 24, and then the distance between 24 and 24.6. Once you have all of those add them together to get your answer.

OpenStudy (anonymous):

Ok, I prefer to stop. There are so many other members here to help you. Do Great!

OpenStudy (anonymous):

But if you need any further explanation, tag me.

OpenStudy (spinex12):

ok

OpenStudy (anonymous):

align your decimal point like this: 24.6 -8.47

OpenStudy (spinex12):

I am so sorry I do not understand

OpenStudy (anonymous):

when subtracting decimals, you have to minus when the decimal point are aligned

OpenStudy (spinex12):

ok now I get it

OpenStudy (spinex12):

finally

OpenStudy (anonymous):

sooo, ya good?

OpenStudy (spinex12):

I think it is b?

OpenStudy (anonymous):

close, it's c

OpenStudy (spinex12):

ok

OpenStudy (anonymous):

yep. see ya

OpenStudy (spinex12):

see ya

OpenStudy (spinex12):

thank you all for your help

OpenStudy (spinex12):

I am closing the question